OpenStudy (itsmichelle29):

The function f(x) = 4(4)x represents the growth of a fly population every year in a remote swamp. Jackie wants to manipulate the formula to an equivalent form that calculates three times a year, not just once a year. Which function is correct for Jackie's purpose, and what is the new growth rate?

OpenStudy (itsmichelle29):

i medal and fan

OpenStudy (itsmichelle29):

f(x) = 4(4)x; growth rate 400% f(x) = 4(4)3x, growth rate 4% f(x) = 4(1.59)3x; growth rate 59% f(x) = 4(1.59)x; growth rate 4%

OpenStudy (triciaal):

instead of x you now have (3x)

OpenStudy (itsmichelle29):

okay so b or c

OpenStudy (itsmichelle29):

is the answer C

OpenStudy (itsmichelle29):

@triciaal

OpenStudy (triciaal):

wait I am reading again 3x 3 times as much or 1/3 the number of growth periods

OpenStudy (itsmichelle29):

s0 D

OpenStudy (triciaal):

sorry don't know

OpenStudy (triciaal):

@surjithayer do you want to help with this?

OpenStudy (itsmichelle29):

can u help me @surjithayer

OpenStudy (itsmichelle29):

Yes or no...